Wo e parts (a) The grade point averages (GPA) for 12 randomly selected college students are shown on the right Complet 0 0 1.6 0.8 40 2.1 1.1 3.5 0.4 2.4 3.3 through (c) below Assume the population is normally distributed (a) Find the sample mearn xRound to two decimal places as needed) (b) Find the sample standard deviation sRound to two decimal places as needed.) (c) Construct a 95% confidence interval for the population mean A 95% confidence interval for the population mean is (Round to two decimal places as needed)
Answer #1

Given Samples of 12 randomly selected are

2.4
1.6
2.1
0.4
3.2
0.8
1.1
2.4
2.9
4
3.5
3.3

a) The mean can be calculared as

2.41.6+ 2.1 0.4+3.2 + 0.8+1.1 2.4 + 2.9+4 + 3.5 +3.3 12 2.31

b) The sample standard deviation calculated as

SD=\sqrt{\frac{\sum (\bar{X}-X)^{2}}{n-1}}

Hence sample standard deviation as 1.14

c). 95 % confidence interval at 95%

Calculation

\bar{X} = 2.31
t = 1.96
sM = √(1.142/12) = 0.33

μ = \bar{X} ± Z(sM)
μ = 2.31 ± 1.96*0.33
μ = 2.31 ± 0.645

Hence 95% CI is [1.67, 2.96].